OpenStudy (vishweshshrimali5):

Can someone explain me please what is sine or cosine rule of triangles............. ?

OpenStudy (vishweshshrimali5):

@satellite73 @mahmit2012

OpenStudy (lgbasallote):

do you mean this thingy \[\frac{a}{\sin A} = \frac{b}{\sin B} = \frac{c}{\sin C}\]

OpenStudy (vishweshshrimali5):

yes

OpenStudy (anonymous):